OpenStudy (anonymous):

he geometric sequence shown can be represented by the function y = 5n, where the domain of the function is 1, 5, 25, 125, ...

OpenStudy (campbell_st):

well the question doesn't make a lot of sense... if the domain is 1 5, 25, .... then the range is y = 5 * 1 , y = 5*5, y = 5*25, y = 5*125,.... 5 25 125 625 ........